40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 8 15:31:48 2013, Jenne, Update, ASC, POP QPD calibration prep 
    Reply  Tue Jul 9 11:41:22 2013, Jenne, Update, ASC, POP QPD calibration attempt 
       Reply  Tue Jul 9 16:08:32 2013, Jenne, Update, ASC, POP QPD calibration attempt 
          Reply  Fri Jul 12 21:23:42 2013, Jenne, Update, ASC, POP QPD calibration attempt ChangeVertMicrometer_July2013.pngChangeSideMicrometer_July2013.png
             Reply  Mon Jul 15 17:16:59 2013, Jenne, Update, ASC, POP QPD calibration attempt ChangeSideMicrometer_July2013_calib2.pngChangeSideMicrometer_July2013_calib1.pngChangeVertMicrometer_July2013_calib1.png
                Reply  Tue Jul 16 01:17:21 2013, Jenne, Update, ASC, POP QPD calibration attempt 
                   Reply  Tue Jul 23 01:30:27 2013, Jenne, Update, ASC, POP QPD analysis 9x
                      Reply  Fri Jul 26 13:39:30 2013, Koji, Update, ASC, POP QPD analysis 
Message ID: 8805     Entry time: Mon Jul 8 15:31:48 2013     Reply to this: 8810
Author: Jenne 
Type: Update 
Category: ASC 
Subject: POP QPD calibration prep 

I am prepping to do the POP QPD calibration, and so have turned off the POP QPD, and put it onto a micrometer stage.  My plan is to (after fixing the ASC servo filters to make the servo AC coupled, rather than DC coupled) lock the PRM-ITMY half cavity, and use that beam to calibrate the QPD.  While this isn't as great as the full PRMI, the PRMI beam moves too much to be useful, unless the ASC servo is engaged.

While on the table, I noticed 2 things:

* In order to place the micrometer, I had to temporarily move the POP55 RFPD (which has not been used in quite a long time).  I think it's just that the panel-mount SMA connector isn't tight to the panel inside, but the RF out SMA cable connector is very loose.  I have moved the POP55 RFPD to the very very south end of the SP table, until someone has time to have a quick look. (I don't want to get too distracted from my current mission, since we haven't put beam onto that PD for at least a year).

* The ITMX oplev beam setup isn't so great.  The last steering mirror before the beam is launched into the vacuum is close to clipping (in yaw... pitch is totally fine), and the steering mirror outside of vacuum to put the beam on the QPD is totally clipping.  The beam is falling off the bottom of this last steering mirror.  Assuming the beam height is okay on all of the input optics and the in-vac table, we need to lower the last steering mirror before the oplev QPD.  My current hypothesis is that by switching which in-vac steering mirror we are using (see Gautam's elog 8758) the new setup has the beam pointing downward a bit.  If the problem is one of the in-vac mirrors, we can't do anything about it until the vent, so for now we can just lower the out of vac mirror.  We should put it back to normal height and fix the oplev setup when we're at atmosphere.

ELOG V3.1.3-